# The Road to Artificial SuperIntelligence: A Comprehensive Survey of Superalignment

### v1

@@ 3 Approaches and Methods
3.1 Weak-to-Strong Generalization (W2SG)
alignment can themselves be exploited by the pri- mary AI they are meant to align, as observed in adversarial vulnerabilities and reward hacking
Figure 3: Weak-to-strong generalization technique. 1 A weak AI system is first trained using instructions and labeled responses. 2 It is then used to generate pseudo-responses for training a strong AI system. Despite the potential noisiness of these pseudo-responses, the strong AI system’s generalization capability (red border) is where essence of scalable oversight comes from, surpassing the performance of the weak AI system.
3.1.1 Definition
3.1.2 Formalization
@@ 3.1.5 Applications
Reasoning One work applies W2SG to complex reasoning tasks on large language models Yang et al. (2024b). It addresses the challenge of enabling strong AI systems to learn from weaker supervisors in reasoning contexts rather than simple classification problems. The proposed progressive learning framework involves two stages—in the first stage, the strong AI system is fine-tuned using a smaller, curated, and more reliable subset of data generated by weak AI systems, combined with in-context learning from the strong AI system itself. In the second stage, preference optimization is applied to learn from the weak AI systems’ errors, enabling the strong AI system to enhance its reasoning capabilities through contrastive sample creation. This approach shows promising results, particularly with larger AI systems (up to 70B parameters) and alignment tasks with direct preference optimization. However, the process of curating reliable training data in the first stage requires careful filtering, which could introduce additional complexity when scaling to larger datasets or different domains.
Vision Guo et al. Guo et al. (2024) applied W2SG to vision foundation models and proposed an adaptive confidence distillation method. This method balances learning from the weaker supervisor while enabling the strong student AI system to rely on its own predictive capabilities when confident. Their approach demonstrates the potential of W2SG in vision tasks. However, a limitation arises from the training setup, where both the weak and strong AI systems are fine-tuned on the original dataset. This overlap makes it challenging to attribute the improvements to W2SG alone.
@@ 3.2.3 Core Concept & Analysis
Irving et al. Irving et al. (2018) introduced the concept of debate, where two AI agents engage in a discussion to persuade a human judge. The judge determines which agent presents the most truthful and useful information. Similar to principles in complexity theory, this framework indicates that the agents are motivated to deliver aligned and truthful arguments. This debate AI system is effective for scalable oversight that is too complex for humans to evaluate directly. However, ASI may exploit the judge’s biases and understanding, possibly misrepresenting human values. Additionally, engaging in debate may require more AI system’s capacity than a direct answer approach, potentially making debate less effective than other training methods. Human judges are also prone to biases, which could lead them to favor arguments that confirm their existing beliefs.
Brown-Cohen et al. Brown-Cohen et al. (2023) enhanced the debate framework by introducing doubly-efficient debate. This debate method enables two polynomial-time provers (i.e., debaters) to compete in convincing a verifier (i.e., judge) of the correctness of a solution while using minimal human judgment. This efficiency in human judgment supports the scalable oversight of AI systems. Meanwhile, the framework provides an efficient method for debate that relies on minimal human judgment. However, there are two key concerns which are the framework assumes the existence of a prover at AGI level, which may not be applicable in practice; and the inherent flaws in human verification reduce the reliability of real-world applications.
Kenton et al. Kenton et al. (2024) investigated scalable oversight using weak LLMs as judges for stronger LLMs. The study examined three protocols—debate, consultancy, and direct question-answering (QA), evaluating them across tasks with varying levels of information asymmetry, including mathematics, coding, and logic. The results showed that debate consistently outperformed consultancy by leveraging adversarial interactions between debaters, enabling weak judges to provide more accurate oversight. However, weak judges’ reliance on debaters and consultants poses challenges that may lead to inaccurate decisions. Furthermore, the study found that both debate and consultancy were less effective than QA protocols where judges directly answered questions or reviewed source articles beforehand, raising questions about the effectiveness of the debate approach.
However, these frameworks assume that debaters generate comprehensible reasoning traces and are vulnerable to the biases or limitations of weak judges. They are also at risk of adversarial exploitation by strong AI systems and may be less reliable due to inherent human limitations, such as bias.
@@ 3.2.4 Enhancement Methods
Kirchner et al. Kirchner et al. (2024) introduced Prover-Verifier Games to improve the legibility of AI system. It involves a “helpful” and “sneaky” prover generating solutions and a verifier evaluating their correctness and iterative refinement. By enabling adversarial interactions, the approach increases the legibility of complex solutions. However, training involves an iterative process between provers and verifiers, requiring multiple rounds and significant computational resources. The reliance on reliable ground truth data limits generalization to domains without labels. Additionally, adversarial provers can exploit verifier weaknesses, and the trade-off between performance and legibility poses challenges to scalability.
Khan et al. Khan et al. (2024) explored the use of more persuasive LLMs in debate setups to enhance the accuracy of weaker judges. Their study demonstrated that optimizing debaters to persuade judges effectively helps non-expert judges better identify truthful answers, particularly in adversarial debates. This optimization improves judge calibration and improves judge performance. However, the results are limited to setups where debaters can provide verified evidence to the judge, which may be problematic if the model relies solely on persuasive output without factual evidence. Additionally, the models are fine-tuned using RLHF for honesty, leaving it unclear whether the debate would remain effective if applied to deceptive models that prioritize performance over truthfulness.
@@ 3.2.5 Applications
Du et al. Du et al. (2023) demonstrated that multi-agent debate enhances reasoning and factuality in language models. This approach involves multiple instances of language models proposing and debating their individual responses and reasoning processes over multiple rounds to reach a common final answer. The method has proven effective in tasks such as mathematical reasoning and factual question-answering. However, the computational complexity increases with the number of debaters involved. Additionally, longer debates can lead to challenges in maintaining coherence and understanding within AI systems.
@@ 3.3.3 Core Concept & Analysis
Bai et al. Bai et al. (2022) introduced the concept of constitutional AI, which aims to train harmless AI systems using a “constitution” of predefined principles. The process includes a supervised learning phase, where the AI system trains with self-generated critiques and revisions, and a reinforcement learning phase that leverages an AI preference model instead of human feedback. Self-critique and self-improvement make Constitutional AI more harmless, while its ability to evaluate critique based on predefined principles enhances transparency. However, reliance on predefined principles may limit adaptability in various environments, as tailored principles are needed. Additionally, using AI feedback instead of human annotations could propagate existing biases within models.
Sharma et al. Sharma et al. (2024) compared two training strategies, SFT and AI feedback, for RL modeling by evaluating their effectiveness, including the capability gap between direct preference training generated by the teacher model and RLAIF training with the AI feedback model. The findings revealed that SFT on high-quality completions can outperform RLAIF. The authors also found that for RLAIF to outperform SFT, a sufficiently strong pre-trained base model is necessary, and there must be a capability mismatch between the teacher used for SFT data collection and the critic used for AI feedback. These results raise concerns that replacing human labelers with AI could lead to suboptimal outcomes.
Yuan et al. Yuan et al. (2024) proposed a a self-rewarding training, where models generate their own feedback during training. This self-improvement paradigm demonstrated improved instruction-following and reward modeling abilities. While promising, this approach demands substantial computational resources and careful calibration.
@@ 3.3.4 Enhancement
Lee et al. Lee et al. (2024) propose Direct-RLAIF (d-RLAIF) by eliminating a separate reward model to train RLAIF. It directly uses feedback from an off-the-shelf LLM during reinforcement learning. This approach addresses key challenges such as RM staleness (i.e., the generated output become increasingly out-of-distribution from the dataset the RM was trained on, leading to suboptimal performance) and the time-consuming of RM training and AI preference labeling. By simplifying the reinforcement learning pipeline, d-RLAIF offers improved efficiency and effectiveness in tasks like summarization and dialogue generation.
@@ 3.3.5 Applications
Yu et al. Yu et al. (2024) introduced RLAIF-V, a framework for aligning multimodal large language models using open-source feedback. Multimodal training is implemented through SFT training with multimodal instructions (e.g., video question-answering tasks). The trained SFT model is used to evaluate preference scores based on detailed descriptions of video clips (which are also self-generated), which are then utilized to train the model as a reward model. Finally, the SFT model undergoes RLAIF training with the reward model using proximal policy optimization (PPO). While RLAIF-V enables multimodal RLAIF training, the method is largely dependent on the quality of the generated content from the base AI model.
Dutta et al. Dutta et al. (2024) applied RLAIF to enhance code generation in lightweight LLMs, focusing on tasks requiring to generate accurate API call. By training on AI-generated feedback from larger models, their framework improved code executability rates significantly, outperforming fine-tuned baselines. However, reliance on larger models for feedback than the base model for training limit the applicability of this approach.
@@ 4 Challenges and Potential Directions
While scalable oversight and alignment paradigms like RLHF, W2SG, and RLAIF have demonstrated notable progress, they still face critical limitations that hinder their ability to address the superalignment problem. Below, we outline the main challenges inherent to the current paradigms and explore potential directions for future work.
@@ 4.2 Potential Directions
Encouraging data diversity can play a pivotal role in fostering creativity and robustness in alignment strategies. Multi-agent collaboration, where agents with different specializations or perspectives interact, has shown promise in enabling emergent behaviors and problem-solving approaches beyond the capabilities of individual models. Competition among agents can further drive improvement, as adversarial settings often reveal diverse argument data that finds alignment weaknesses and suggests novel pathways for improvement. However, diversity-based approaches require careful curation and filtering to ensure that resulting behaviors are beneficial and do not inadvertently reinforce undesirable traits.
@@ 5 Conclusion
In this survey, we explored superalignment, its challenges, and scalable oversight techniques. We analyzed current paradigms, and practical methodologies, and their potential and limitations. By identifying critical challenges and proposing future directions, we aim to bridge current alignment techniques with the goal of achieving safe and aligned ASI. Addressing these challenges is vital for guiding AI systems toward not only avoiding harm but benefiting humanity—a crucial step toward a future where humans and AI collaborate to solve complex global problems.

### v3

@@ 3 Approaches and Methods
3.1 Weak-to-Strong Generalization (W2SG)
Figure 2: Weak-to-strong generalization technique. 1 A weak AI system is first trained using instructions and labeled responses. 2 It is then used to generate pseudo-responses for training a strong AI system. Despite the potential noisiness of these pseudo-responses, the strong AI system’s generalization capability (red border) is where essence of scalable oversight comes from, surpassing the performance of the weak AI one.
3.1.1 Definition
3.1.2 Formalization
@@ 3.1.5 Applications
Reasoning
One work applies W2SG to complex reasoning tasks on large language models Yang et al. (2024b). It addresses the challenge of enabling strong AI systems to learn from weaker supervisors in reasoning contexts rather than simple classification problems. The proposed progressive learning framework involves two stages—in the first stage, the strong AI system is fine-tuned using a smaller, curated, and more reliable subset of data generated by weak AI systems, combined with in-context learning from the strong AI system itself. In the second stage, preference optimization is applied to learn from the weak AI systems’ errors, enabling the strong AI system to enhance its reasoning capabilities through contrastive sample creation. This approach shows promising results, particularly with larger AI systems (up to 70B parameters) and alignment tasks with direct preference optimization. However, the process of curating reliable training data in the first stage requires careful filtering, which could introduce additional complexity when scaling to larger datasets or different domains.
Vision
Guo et al. (2024) applied W2SG to vision foundation models and proposed an adaptive confidence distillation method. This method balances learning from the weaker supervisor while enabling the strong student AI system to rely on its own predictive capabilities when confident. Their approach demonstrates the potential of W2SG in vision tasks. However, a limitation arises from the training setup, where both the weak and strong AI systems are fine-tuned on the original dataset. This overlap makes it challenging to attribute the improvements to W2SG alone.
@@ 3.2.3 Core Concept & Analysis
Irving et al. (2018) introduced the concept of debate, where two AI agents engage in a discussion to persuade a human judge. The judge determines which agent presents the most truthful and useful information. Similar to principles in complexity theory, this framework indicates that the agents are motivated to deliver aligned and truthful arguments. This debate AI system is effective for scalable oversight that is too complex for humans to evaluate directly. However, ASI may exploit the judge’s biases and understanding, possibly misrepresenting human values. Additionally, engaging in debate may require more AI system’s capacity than a direct answer approach, potentially making debate less effective than other training methods. Human judges are also prone to biases, which could lead them to favor arguments that confirm their existing beliefs.
Brown-Cohen et al. (2023) enhanced the debate framework by introducing doubly-efficient debate. This debate method enables two polynomial-time provers (i.e., debaters) to compete in convincing a verifier (i.e., judge) of the correctness of a solution while using minimal human judgment. This efficiency in human judgment supports the scalable oversight of AI systems. Meanwhile, the framework provides an efficient method for debate that relies on minimal human judgment. However, there are two key concerns which are the framework assumes the existence of a prover at AGI level, which may not be applicable in practice; and the inherent flaws in human verification reduce the reliability of real-world applications.
Kenton et al. (2024) investigated scalable oversight using weak LLMs as judges for stronger LLMs. The study examined three protocols—debate, consultancy, and direct question-answering (QA), evaluating them across tasks with varying levels of information asymmetry, including mathematics, coding, and logic. The results showed that debate consistently outperformed consultancy by leveraging adversarial interactions between debaters, enabling weak judges to provide more accurate oversight. However, weak judges’ reliance on debaters and consultants poses challenges that may lead to inaccurate decisions. Furthermore, the study found that both debate and consultancy were less effective than QA protocols where judges directly answered questions or reviewed source articles beforehand, raising questions about the effectiveness of the debate approach.
However, these frameworks assume that debaters generate comprehensible reasoning traces and are vulnerable to the biases or limitations of weak judges. They are also at risk of adversarial exploitation by strong AI systems and may be less reliable due to inherent human limitations.
@@ 3.2.4 Enhancement Methods
Kirchner et al. (2024) introduced Prover-Verifier Games to improve the legibility of AI system. It involves a “helpful” and “sneaky” prover generating solutions and a verifier evaluating their correctness and iterative refinement. By enabling adversarial interactions, the approach increases the legibility of complex solutions. However, training involves an iterative process between provers and verifiers, requiring multiple rounds and significant computational resources. The reliance on reliable ground truth data limits generalization to domains without labels. Additionally, adversarial provers can exploit verifier weaknesses, and the trade-off between performance and legibility poses challenges to scalability.
Khan et al. (2024) explored the use of more persuasive LLMs in debate setups to enhance the accuracy of weaker judges. Their study demonstrated that optimizing debaters to persuade judges effectively helps non-expert judges better identify truthful answers, particularly in adversarial debates. This optimization improves judge calibration and improves judge performance. However, the results are limited to setups where debaters can provide verified evidence to the judge, which may be problematic if the model relies solely on persuasive output without factual evidence. Additionally, the models are fine-tuned using RLHF for honesty, leaving it unclear whether the debate would remain effective if applied to deceptive models that prioritize performance over truthfulness.
3.2.5 Evaluation
The SCALEEVAL leverages multi-agent debates to establish scalable and reliable evaluation benchmarks for LLMs across diverse scenarios, such as brainstorming Chern et al. (2024). By enabling agents to critique and iteratively discuss responses, it minimizes the reliance on extensive human annotations while maintaining alignment with expert judgments. However, its dependence on agent agreement and limited human involvement can pose challenges, particularly in edge cases where consensus is difficult to achieve.
@@ 3.2.6 Applications
Du et al. (2023) demonstrated that multi-agent debate enhances reasoning and factuality in language models. This approach involves multiple instances of language models proposing and debating their individual responses and reasoning processes over multiple rounds to reach a common final answer. The method has proven effective in tasks such as mathematical reasoning and factual question-answering. However, the computational complexity increases with the number of debaters involved. Additionally, longer debates can lead to challenges in maintaining coherence and understanding within AI systems.
@@ 3.3.3 Core Concept & Analysis
Bai et al. (2022) introduced the concept of constitutional AI, which aims to train harmless AI systems using a “constitution” of predefined principles. The process includes a supervised learning phase, where the AI system trains with self-generated critiques and revisions, and a reinforcement learning phase that leverages an AI preference model instead of human feedback. Self-critique and self-improvement make Constitutional AI more harmless, while its ability to evaluate critique based on predefined principles enhances transparency. However, reliance on predefined principles may limit adaptability in various environments, as tailored principles are needed. Additionally, using AI feedback instead of human annotations could propagate existing biases within models.
Sharma et al. (2024) compared two training strategies, SFT and AI feedback, for RL modeling by evaluating their effectiveness, including the capability gap between direct preference training generated by the teacher model and RLAIF training with the AI feedback model. The findings revealed that SFT on high-quality completions can outperform RLAIF. The authors also found that for RLAIF to outperform SFT, a sufficiently strong pre-trained base model is necessary, and there must be a capability mismatch between the teacher used for SFT data collection and the critic used for AI feedback. These results raise concerns that replacing human labelers with AI could lead to suboptimal outcomes.
Yuan et al. (2024) proposed a a self-rewarding training, where models generate their own feedback during training. This self-improvement paradigm demonstrated improved instruction-following and reward modeling abilities. While promising, this approach demands substantial computational resources and careful calibration.
@@ 3.3.4 Enhancement
Lee et al. (2024) propose Direct-RLAIF (d-RLAIF) by eliminating a separate reward model to train RLAIF. It directly uses feedback from an off-the-shelf LLM during reinforcement learning. This approach addresses key challenges such as RM staleness (i.e., the generated output become increasingly out-of-distribution from the dataset the RM was trained on, leading to suboptimal performance) and the time-consuming of RM training and AI preference labeling. By simplifying the reinforcement learning pipeline, d-RLAIF offers improved efficiency and effectiveness in tasks like summarization and dialogue generation.
@@ 3.3.5 Applications
Yu et al. (2024) introduced RLAIF-V, a framework for aligning multimodal large language models using open-source feedback. Multimodal training is implemented through SFT training with multimodal instructions (e.g., video question-answering tasks). The trained SFT model is used to evaluate preference scores based on detailed descriptions of video clips (which are also self-generated), which are then utilized to train the model as a reward model. Finally, the SFT model undergoes RLAIF training with the reward model using proximal policy optimization (PPO). While RLAIF-V enables multimodal RLAIF training, the method is largely dependent on the quality of the generated content from the base AI model.
Dutta et al. (2024) applied RLAIF to enhance code generation in lightweight LLMs, focusing on tasks requiring to generate accurate API call. By training on AI-generated feedback from larger models, their framework improved code executability rates significantly, outperforming fine-tuned baselines. However, reliance on larger models for feedback than the base model for training limit the applicability of this approach.
@@ 4 Challenges and Potential Directions
While scalable oversight and alignment paradigms like RLHF, W2SG, and RLAIF have demonstrated notable progress, they still face critical limitations that hinder their ability to address the superalignment problem. In this section, we outline the main challenges inherent to the current paradigms and explore potential directions for future work.
@@ 4.2 Potential Directions
Encouraging data diversity can play a pivotal role in fostering creativity and robustness in alignment strategies. Multi-agent collaboration, where agents with different specializations or perspectives interact, has shown promise in enabling emergent behaviors and problem-solving approaches beyond the capabilities of individual models. Competition among agents can further drive improvement, as adversarial settings often reveal diverse argument data that finds alignment weaknesses and suggests novel pathways for improvement. However, diversity-based approaches require careful filtering to ensure that resulting behaviors are beneficial and do not inadvertently reinforce undesirable traits.
@@ 5 Conclusion
In this survey, we explored superalignment, its challenges, and scalable oversight techniques. We analyzed current paradigms, and practical methodologies, and their potential and limitations. By identifying critical challenges and proposing future directions, we aim to bridge current alignment techniques with the goal of achieving safe and aligned AI, preparing for future more advanced AI. Addressing these challenges is vital for guiding AI systems toward not only avoiding harm but benefiting humanity—a crucial step toward a future where humans and AI collaborate to solve complex global real-world problems.